OpenStudy (anonymous):

if cos x = 3/8 what is sin x?

OpenStudy (whpalmer4):

Use the trig identity \[\cos^2{x} + \sin^2{x} = 1\]

OpenStudy (whpalmer4):

Remember that sin, cos are defined in terms of that unit circle, so the hypotenuse is always 1, and the sides of the right triangle are sin and cos

OpenStudy (anonymous):

Alright, thanks.

OpenStudy (anonymous):

Though it's been like 2 years since I've done this...

OpenStudy (anonymous):

I ended up with like x coming to be .002454334215773

OpenStudy (anonymous):

and plugged it into the trig identity equation and got very close to one, so I assume that's the correct answer.

OpenStudy (whpalmer4):

Let's work it out. \[\cos{x} = 3/8\]\[\cos^2x+\sin^2x = 1\]\[(3/8)^2+\sin^2x = 1\]\[\sin^2x = \frac{64}{64}-\frac{9}{64}\]\[\sin x=\frac{\sqrt{55}}{8} \approx 0.927 \]

OpenStudy (whpalmer4):

If you did this with a calculator, is your calculator expecting degrees, or radians for trig functions? The identity applies to radians, not degrees.

OpenStudy (whpalmer4):

@modestgengar

OpenStudy (anonymous):

I forgot to divide by 8...that's where I went wrong.